# RobustVLA: Robustness-Aware Reinforcement Post-Training for Vision-Language-Action Models

### v1

@@ 4.2 Model online regularization optimization objective
On the other hand, Theorem 2 states that the return gap can be bounded in terms of the worst-case deviation $\delta_{t}=\sup_{s\in\mathcal{S}}\|\pi_{\theta}(\cdot|s)-\pi_{\theta^{-}}(\cdot|s)\|_{\infty}$, which captures the maximum discrepancy between the new and old models across the entire state space. Although this characterization is theoretically precise, computing or constraining $\delta_{t}$ directly is infeasible in practice, especially in continuous action domains. To operationalize this idea, we introduce easy-to-handle alternatives. First, the PPO objective already constrains the average distributional shift through a KL divergence penalty: $\mathbb{E}_{s\sim\mathcal{D}}\Big[D_{\mathrm{KL}}\big(\pi_{\theta}(\cdot|s)\,\|\,\pi_{\theta^{-}}(\cdot|s)\big)\Big]\leq\epsilon.$ By Pinsker’s inequality, this implies $\|\pi_{\theta}(\cdot|s)-\pi_{\theta^{-}}(\cdot|s)\|_{1}\leq\sqrt{2D_{\mathrm{KL}}\big(\pi_{\theta}(\cdot|s)\|\pi_{\theta^{-}}(\cdot|s)\big)},$ thus providing an average-case upper bound on model deviation. However, the KL penalty tends to be sensitive to variance changes and does not necessarily prevent large shifts in the mean action outputs. To complement this, we further introduce a smoothness regularizer that explicitly penalizes deviations in the model means:
@@
where $\mu_{\theta}(s)$ denotes the mean action. This regularizer directly bounds the displacement of the mean actions, which in turn controls a surrogate of $\delta_{t}$ in expectation. Intuitively, while the KL constraint ensures distribution-level stability, the $\mathcal{R}_{\text{Smooth}}$ enforces gradual updates in the action means, thereby preventing erratic model shifts that may otherwise amplify error accumulation. This dual mechanism—average KL regularization together with mean-level $L_{2}$ smoothing—serves as a practical surrogate for bounding $\delta_{t}$, thus aligning the algorithmic implementation with the theoretical insights of Theorem 2. Therefore, taking these regularization terms into consideration, we obtain a robust online RL optimization objective:
@@ 5.1 Performance Evaluation under Environmental Perturbations
To comprehensively evaluate the performance of the RobustVLA, we construct a robust testing benchmark based on the LIBERO (Liu et al., 2023) simulation platform that incorporates environmental uncertainty and perturbations, as shown in Fig. 2. The original LIBERO benchmark is a long-horizon robotic learning suite consisting of Objects, Long, Spatial, and Goal. We impose five observation perturbations and three action perturbations during the online interaction between the VLA model and the environment. The five observation perturbations are Image Shift, Image Rotation, Color Jitter, Image Occlusions, and Image Erasing. The three action perturbations are zero-mean Gaussian action noise with standard deviations of 0.1, 0.2, and 0.3. Details of the observation perturbations are in Appendix A.4. For model evaluation, we follow the common scheme of previous work (Tan et al., 2025a), where a single VLA model is deployed to all tasks in the suite, and the deployment is performed on 50 held-out test contexts for each task.
@@
Our proposed algorithm comes in two versions: one without curriculum learning (RobustVLA) and one with curriculum learning (RobustVLA-C). For a thorough comparison, three categories of state-of-the-art baseline models are considered. The first category is Offline Imitation Learning (Offline IL) models, which are trained solely on static expert datasets. These include: 1) GEVRM (Zhang et al., 2025a): This model uses prototype contrastive learning to resist observation perturbations. 2) OpenVLA (Kim et al., 2024), which represents robot actions as discrete token sequences and predicts them one by one in sequence. 3) OpenVLA-OFT (Kim et al., 2025), which follows the OpenVLA modeling philosophy and further proposes improvements such as parallel decoding. 4) $\pi_{0}$ (Black et al., 2024), a VLA flow model that effectively models multi-modal distributions. The second category is Offline Reinforcement Learning (Offline RL) models, which apply RL algorithms to offline datasets to more fully capture the data quality distribution. These include: 5) RWR (Peters & Schaal, 2007), which optimizes the model by performing reward-weighted regression on samples. 6) ReinboT (Zhang et al., 2025c), which guides VLA action generation by predicting densely maximized future returns. 7) ARFM (Zhang et al., 2025b), a fine-tuning algorithm that adaptively adjusts the weights of a flow-matching RL loss. The final category is Online Reinforcement Learning (Online RL) models, which fine-tune a pre-trained VLA model by autonomously interacting with the environment. This includes 8) RIPT-VLA (Tan et al., 2025a), a critic-free fine-tuning algorithm that uses only sparse binary success rewards. The implementation details of our method and the reproduction of the baseline algorithm are shown in Appendix A.5.
@@ 5.2 Transfer Learning Performance Comparison
In contrast, on the ”open drawer” and ”put bowl” tasks, RobustVLA achieves improvements of $8.0\%$ and $16.0\%$ over direct zero-shot transfer, respectively. This demonstrates that the proposed RobustVLA can significantly withstand environmental perturbations and adopt more robust and stable behaviors after only a few environment rollouts, thus exhibiting superior transfer learning capabilities. More results in Appendix Fig. 5.
@@ 5.3 Ablation Analysis
Observation Representation Visualization. To investigate how the proposed penalty objective affects the observation representation of the VLA model, we performed a T-SNE (Maaten & Hinton, 2008) visualization analysis in the same joint perturbation setting, as shown in Fig. 4(b-c). The result shows that observation representations corresponding to the RobustVLA remain largely unchanged despite varying degrees of environmental perturbation. This demonstrates its robustness to environmental uncertainty, enabling robust task performance. In contrast, the baseline RIPT-VLA exhibits overly separated clusters in the observation representations of some downstream tasks. This indicates a deviation from the desired trajectory for task performance, leading to a lack of robustness and reliability in decision-making.
@@ A.1 More related Work
Recent advances in foundation models have extended large-scale pretraining beyond language and vision to embodied agents, giving rise to Vision-Language-Action (VLA) models. The initial wave of this paradigm was marked by two key approaches: extending robotics-native architectures to internet scale, exemplified by RT-1 (Brohan et al., 2022) and RT-2 (Zitkovich et al., 2023), and injecting multimodal sensor data directly into large language models, with PaLM-E (Driess et al., 2023) representing a landmark embodied language model. Subsequent efforts explored improved architectures and scaling laws, with models such as Octo (Team et al., 2024) and OpenVLA (Kim et al., 2024) establishing powerful, open-source benchmarks. Architectural exploration also led to new policy representations, exemplified by the use of diffusion transformers in Dita (Hou et al., 2025) for scaling generalist policies. More recent work has intensified the focus on grounding VLAs in high-quality, robotics-specific data, including the aggregation of diverse datasets via projects like RT-X (O’Neill et al., 2024), GR-3 (Cheang et al., 2025), and pioneering data collection methods such as Mobile ALOHA (Fu et al., 2024). Meanwhile, to advance real-world generalization, models including H-RDT (Bi et al., 2025), $\pi_{0}$ (Black et al., 2024) and $\pi_{0.5}$ (Intelligence et al., 2025) have been developed to improve embodied reasoning and address open-world, knowledge-driven tasks. Collectively, these studies position VLA models as a core foundation for general-purpose embodied intelligence. Moreover, in the field of robust vision-language manipulation, previous work, GEVRM (Zhang et al., 2025a), evaluates environmental perturbations by simulating responses and optimizing them through prototype contrastive learning. However, GEVRM is inherently a hierarchical imitation learning framework, and the distribution shift of its visual planning objectives during training and testing can lead to performance bottlenecks. Other work (Hancock et al., 2025) has proposed a runtime intervention scheme that identifies model-sensitive input regions and modifies task-irrelevant regions to reduce model sensitivity, thereby improving visual robustness. However, this approach relies heavily on off-the-shelf VLMs and segmentation models, and its primary limitation is accurately distinguishing objects from background clutter. In contrast to these works, our work focuses on online RL post-training to further adapt and coordinate the VLA model with downstream tasks.
@@ A.2 Further analysis of model robustness
Under the setting of Theorem 1, suppose the imitation VLA model is optimal ($\epsilon_{\text{offline}}=0$) and the dynamics are contractive ($L_{f}<1$). Then the return gap satisfies:
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(H)\cdot L_{r}L_{f}\cdot\lambda\epsilon_{s}.$
Corollary 1 shows that the compounding factor $\mathcal{O}(L_{f}^{H})$ collapses to a linear dependence on the horizon $H$, yielding a return gap on the order of $H\cdot\lambda\epsilon_{s}$. What this implies is that observation noise no longer causes exponential amplification. Instead, the performance degradation grows linearly with both the horizon and the product $\lambda\epsilon_{s}$. This highlights a regime where Jacobian regularization has direct leverage -- by shrinking $\lambda$, the sensitivity to $\epsilon_{s}$ is proportionally reduced, ensuring that perception errors do not destabilize the system even in long rollouts. In practice, this means that in sufficiently stable environments, robustness to noisy observations is achievable with only modest regularization strength.
@@
Under the setting of Theorem 3 but with discounted return $J(\pi)=\mathbb{E}\Big[\sum_{t\geq 1}\gamma^{t-1}r(s_{t},a_{t})\Big]$ for some discount factor $\gamma\in(0,1)$. The dynamics Lipschitz constant and discount satisfy $\gamma L_{f}<1$.The per-step driving term $\zeta_{t}:=\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}+\sum_{i=1}^{t}\delta_{i}+\|\xi_{t}\|$ is uniformly bounded: $\sup_{t}\zeta_{t}\leq C<\infty$. The model deviation sequence $\{\epsilon_{t}\}$ satisfies $\sup_{t}\epsilon_{t}\leq E<\infty$. The per-step update bound $\|\pi_{t}-\pi_{t-1}\|_{\infty}\leq\delta$ with $\delta=o(1/H)$ as $H\to\infty$ (so that, with a good initialization, $\sup_{t}\epsilon_{t}=O(1)$). Then the return gap is bounded independently of horizon $H$:
$J(\pi^{*})-J(\pi)=\mathcal{O}(1).$
Corollary 2 shows that the nonstationarity term $\sum_{t=1}^{H}\delta_{t}$ remains bounded by a constant independent of $H$, and the return gap simplifies to $\mathcal{O}(1)$. The implication is that with sufficiently strong smoothness regularization, the long-horizon error drift is effectively neutralized: even as $H\to\infty$, the cumulative deviation does not grow. This shows that robustness against action perturbations is not merely about reducing the variance term $\sigma\sqrt{d}$, but also about ensuring that the model evolution is slow enough that compounding drift vanishes asymptotically. Practically, this provides a precise criterion: as long as $\delta$ decays faster than $1/H$, one can guarantee stable long-horizon adaptation without robustness loss.
@@ A.3 Proof of robustness theories
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\lambda\epsilon_{s}\right).$
@@
According to Lemma 1, expanding the recursive formula of $d_{0}=0$ yields:
$d_{t}\leq L_{f}(\lambda\epsilon_{s}+\epsilon_{\text{offline}})\cdot\sum_{i=0}^{t-1}L_{f}^{i}=(\lambda\epsilon_{s}+\epsilon_{\text{offline}})\cdot L_{f}\cdot\frac{L_{f}^{t}-1}{L_{f}-1},$
which is $\mathcal{O}(L_{f}^{t})$.
At each step:
$|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\leq L_{r}\cdot(d_{t}+\|a_{t}-a_{t}^{*}\|)\leq L_{r}\cdot(d_{t}+\lambda\epsilon_{s}+\epsilon_{\text{offline}}).$
Summing over $H$ steps:
$J(\pi^{*})-J(\pi)\leq\sum_{t=1}^{H}L_{r}\cdot(d_{t}+\lambda\epsilon_{s}+\epsilon_{\text{offline}}),$
and since $d_{t}=\mathcal{O}(L_{f}^{t}(\lambda\epsilon_{s}+\epsilon_{\text{offline}}))$, we obtain:
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot(\lambda\epsilon_{s}+\epsilon_{\text{offline}}).$
@@
Let $d_{t}=\|s_{t}-s_{t}^{*}\|$. Then under Lipschitz dynamics:
$d_{t+1}\leq L_{f}\cdot(d_{t}+\epsilon_{t}+\|\xi_{t}\|),$
where $\epsilon_{t}=\|\pi_{t}-\pi^{*}\|\leq\epsilon_{\text{offline}}+\sum_{i=1}^{t}\delta_{i}$.
@@
Then:
$d_{t+1}=\|f(s_{t},a_{t})-f(s_{t}^{*},a_{t}^{*})\|\leq L_{f}(\|s_{t}-s_{t}^{*}\|+\|a_{t}-a_{t}^{*}\|),$
so:
$d_{t+1}\leq L_{f}\cdot(d_{t}+\epsilon_{t}+\|\xi_{t}\|).$
@@
Assume actions are perturbed as $a_{t}=\pi_{t}(s_{t})+\xi_{t}$ with $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$, and VLA models satisfy $\|\pi_{t}-\pi_{t-1}\|_{\infty}\leq\delta_{t}$. Then:
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\sum_{t=1}^{H}\delta_{t}+\sigma\sqrt{d}\right).$
@@
From the smooth update constraint:
$\epsilon_{t}=\|\pi_{t}-\pi^{*}\|\leq\|\pi_{init}-\pi^{*}\|+\sum_{i=1}^{t}\|\pi_{i}-\pi_{i-1}\|\leq\epsilon_{\text{offline}}+\sum_{i=1}^{t}\delta_{i}.$
Define:
$\beta_{t}:=L_{f}(\epsilon_{\text{offline}}+\sum_{i=1}^{t}\delta_{i}+\|\xi_{t}\|),\quad d_{0}=0.$
According to Lemma 2, we have:
$d_{t}\leq\sum_{i=0}^{t-1}L_{f}^{t-1-i}\cdot\beta_{i}.$
This implies:
$d_{t}\leq L_{f}^{t-1}\cdot\max_{i\leq t}\beta_{i}\cdot t\leq\mathcal{O}(L_{f}^{t})\cdot\left(\epsilon_{\text{offline}}+\sum_{i=1}^{t}\delta_{i}+\|\xi_{t}\|\right).$
At each time step, reward error:
$|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\leq L_{r}(d_{t}+\|a_{t}-a_{t}^{*}\|)\leq L_{r}(d_{t}+\epsilon_{t}+\|\xi_{t}\|).$
Summing over $H$ steps:
$J(\pi^{*})-J(\pi)\leq\sum_{t=1}^{H}L_{r}\cdot(d_{t}+\epsilon_{t}+\|\xi_{t}\|).$
Further according to Lemma 3, we obtain:
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\sum_{t=1}^{H}\delta_{t}+\sigma\sqrt{d}\right).$

### v2

@@ 4.2 Model online regularization optimization objective
On the other hand, Theorem 2 states that the return gap can be bounded in terms of the worst-case deviation ${\delta_{i}}=\sup_{s\in\mathcal{S}}\|\pi_{\theta}(\cdot|s)-\pi_{\theta^{-}}(\cdot|s)\|_{\infty}$, which captures the maximum discrepancy between the new and old models across the entire state space. Although this characterization is theoretically precise, computing or constraining $\delta_{i}$} directly is infeasible in practice, especially in continuous action domains. To operationalize this idea, we introduce easy-to-handle alternatives. First, the PPO objective already constrains the average distributional shift through a KL divergence penalty: $\mathbb{E}_{s\sim\mathcal{D}}\Big[D_{\mathrm{KL}}\big(\pi_{\theta}(\cdot|s)\,\|\,\pi_{\theta^{-}}(\cdot|s)\big)\Big]\leq\epsilon.$ By Pinsker’s inequality, this implies $\|\pi_{\theta}(\cdot|s)-\pi_{\theta^{-}}(\cdot|s)\|_{1}\leq\sqrt{2D_{\mathrm{KL}}\big(\pi_{\theta}(\cdot|s)\|\pi_{\theta^{-}}(\cdot|s)\big)},$ thus providing an average-case upper bound on model deviation. However, the KL penalty tends to be sensitive to variance changes and does not necessarily prevent large shifts in the mean action outputs. To complement this, we further introduce a smoothness regularizer that explicitly penalizes deviations in the model means:
@@
where $\mu_{\theta}(s)$ denotes the mean action. This regularizer directly bounds the displacement of the mean actions, which in turn controls a surrogate of $\delta_{i}$} in expectation. Intuitively, while the KL constraint ensures distribution-level stability, the $\mathcal{R}_{\text{Smooth}}$ enforces gradual updates in the action means, thereby preventing erratic model shifts that may otherwise amplify error accumulation. This dual mechanism -- average KL regularization together with mean-level $L_{2}$ smoothing -- serves as a practical surrogate for bounding $\delta_{i}$}, thus aligning the algorithmic implementation with the theoretical insights of Theorem 2. Therefore, taking these regularization terms into consideration, we obtain a robust online RL optimization objective:
@@ 5.1 Performance Evaluation under Environmental Perturbations
To comprehensively evaluate the performance of the RobustVLA, we construct a robust testing benchmark based on the LIBERO (liu2023libero) simulation platform that incorporates environmental uncertainty and perturbations, as shown in Fig. 2. The original LIBERO benchmark is a long-horizon robotic learning suite consisting of Objects, Long, Spatial, and Goal. We impose five observation perturbations and three action perturbations during the online interaction between the VLA model and the environment. The five observation perturbations are Image Shift, Image Rotation, Color Jitter, Image Occlusions, and Image Erasing. The three action perturbations are zero-mean Gaussian action noise with standard deviations of 0.1, 0.2, and 0.3. Details of the observation perturbations are in Appendix A.4. For model evaluation, we follow the common scheme of previous work (tan2025interactive), where a single VLA model is deployed to all tasks in the suite, and the deployment is performed on 50 held-out test contexts for each task.
@@
Our proposed algorithm comes in two versions: one without curriculum learning (RobustVLA) and one with curriculum learning (RobustVLA-C). For a thorough comparison, three categories of state-of-the-art baseline models are considered. The first category is Offline Imitation Learning (Offline IL) models, which are trained solely on static expert datasets. These include: 1) GEVRM (zhang2025gevrm): This model uses prototype contrastive learning to resist observation perturbations. 2) OpenVLA (kim2024openvla), which represents robot actions as discrete token sequences and predicts them one by one in sequence. 3) OpenVLA-OFT (kim2025fine), which follows the OpenVLA modeling philosophy and further proposes improvements such as parallel decoding. 4) $\pi_{0}$ (black2024pi_0), a VLA flow model that effectively models multi-modal distributions. The second category is Offline Reinforcement Learning (Offline RL) models, which apply RL algorithms to offline datasets to more fully capture the data quality distribution. These include: 5) RWR (peters2007reinforcement), which optimizes the model by performing reward-weighted regression on samples. 6) ReinboT (zhang2025reinbot), which guides VLA action generation by predicting densely maximized future returns. 7) ARFM (zhang2025balancingsignalvarianceadaptive), a fine-tuning algorithm that adaptively adjusts the weights of a flow-matching RL loss. The final category is Online Reinforcement Learning (Online RL) models, which fine-tune a pre-trained VLA model by autonomously interacting with the environment. This includes 8) RIPT-VLA (tan2025interactive), a critic-free fine-tuning algorithm that uses only sparse binary success rewards. The implementation details of our method and the reproduction of the baseline algorithm are shown in Appendix A.5.
@@ 5.2 Transfer Learning Performance Comparison
In contrast, on the open drawer and put bowl tasks, RobustVLA achieves improvements of $8.0\%$ and $16.0\%$ over direct zero-shot transfer, respectively. This demonstrates that the proposed RobustVLA can significantly withstand environmental perturbations and adopt more robust and stable behaviors after only a few environment rollouts, thus exhibiting superior transfer learning capabilities. More results in Appendix Fig. 6.
@@ 5.3 Ablation Analysis
Observation Representation Visualization. To investigate how the proposed penalty objective affects the observation representation of the VLA model, we performed a T-SNE (maaten2008visualizing) visualization analysis in the same joint perturbation setting, as shown in Fig. 4(b-c). The result shows that observation representations corresponding to the RobustVLA remain largely unchanged despite varying degrees of environmental perturbation. This demonstrates its robustness to environmental uncertainty, enabling robust task performance. In contrast, the baseline RIPT-VLA exhibits overly separated clusters in the observation representations of some downstream tasks. This indicates a deviation from the desired trajectory for task performance, leading to a lack of robustness and reliability in decision-making.
Figure 5: (a) Action distribution during model inference under four noise conditions. (b) Comparison of action distribution between baseline RIPT-VLA and the proposed RobustVLA.
Action Distribution Visualization. We visualized the action distribution to explore how various noises affect model execution and to compare the proposed RobustVLA with the baseline RIPT-VLA. Specifically, we recorded $100$ action trajectories during RIPT-VLA inference in LIBERO Spatial. Four settings were utilized: 1) no observation noise or action noise; 2) only image rotation perturbation; 3) only action noise (noise level $0.1$); and 4) both image rotation perturbation and action noise. The visualization results are shown in Fig. 5(a) and Appendix Fig. 7. The results show that the action trajectories exhibit diverse distribution patterns and coverage under different noise conditions. This indicates that observation perturbation and action perturbation have different emphases and degrees of influence on robot manipulation behavior, especially when both perturbations are present. Furthermore, the action distribution comparison between the baseline RIPT-VLA and the proposed RobustVLA is shown in Fig. 5(b) and Appendix Fig. 8. The results show that under joint perturbation, the action distribution of RobustVLA is closer to the action distribution under noise-free conditions compared to RIPT-VLA. This demonstrates that RobustVLA can adaptively adjust its behavior, enabling the robot to better resist environmental disturbances and thus produce more robust and stable decision-making actions.
@@ A.1 More related Work
Recent advances in foundation models have extended large-scale pretraining beyond language and vision to embodied agents, giving rise to Vision-Language-Action (VLA) models. The initial wave of this paradigm was marked by two key approaches: extending robotics-native architectures to internet scale, exemplified by RT-1 (brohan2022rt) and RT-2 (zitkovich2023rt), and injecting multimodal sensor data directly into large language models, with PaLM-E (driess2023palm) representing a landmark embodied language model. Subsequent efforts explored improved architectures and scaling laws, with models such as Octo (team2024octo) and OpenVLA (kim2024openvla) establishing powerful, open-source benchmarks. Architectural exploration also led to new policy representations, exemplified by the use of diffusion transformers in Dita (hou2025dita) for scaling generalist policies. More recent work has intensified the focus on grounding VLAs in high-quality, robotics-specific data, including the aggregation of diverse datasets via projects like RT-X (o2024open), GR-3 (cheang2025gr), and pioneering data collection methods such as Mobile ALOHA (fu2024mobile). Meanwhile, to advance real-world generalization, models including H-RDT (bi2025h), $\pi_{0}$ (black2024pi_0) and $\pi_{0.5}$ (intelligence2025pi_) have been developed to improve embodied reasoning and address open-world, knowledge-driven tasks. Collectively, these studies position VLA models as a core foundation for general-purpose embodied intelligence. Moreover, in the field of robust vision-language manipulation, previous work, GEVRM (zhang2025gevrm), evaluates environmental perturbations by simulating responses and optimizing them through prototype contrastive learning. However, GEVRM is inherently a hierarchical imitation learning framework, and the distribution shift of its visual planning objectives during training and testing can lead to performance bottlenecks. Other work (hancock2025run) has proposed a runtime intervention scheme that identifies model-sensitive input regions and modifies task-irrelevant regions to reduce model sensitivity, thereby improving visual robustness. However, this approach relies heavily on off-the-shelf VLMs and segmentation models, and its primary limitation is accurately distinguishing objects from background clutter. In contrast to these works, our work focuses on online RL post-training to further adapt and coordinate the VLA model with downstream tasks.
@@ A.2 Further analysis of model robustness
Under the assumptions of Theorem 1 but with $\epsilon_{\mathrm{offline}}=0$ (the imitation model is perfect) and contractive dynamics $L_{f}<1$, the expected return gap satisfies
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\;\leq\;H\cdot L_{r}\cdot\frac{1}{1-L_{f}}\cdot\lambda\epsilon_{s}.$}
Corollary 1 shows that the compounding factor $\mathcal{O}(L_{f}^{H})$ collapses to a linear dependence on the rollout horizon $H$, yielding a return gap on the order of $H\cdot\lambda\epsilon_{s}$. What this implies is that observation noise no longer causes exponential amplification. This highlights a regime where Jacobian regularization has direct leverage -- by shrinking $\lambda$, the sensitivity to $\epsilon_{s}$ is proportionally reduced, ensuring that perception errors do not destabilize the system even in long rollouts. In practice, this means that in sufficiently stable environments, robustness to noisy observations is achievable with only modest regularization strength.
@@
Under the setting of Theorem 3 but with discounted return $J(\pi)=\mathbb{E}\Big[\sum_{t\geq 1}\gamma^{t-1}r(s_{t},a_{t})\Big]$, for some discount factor $\gamma\in(0,1)$ with contractive dynamics $L_{f}\in(0,1)$. Assume the per-step driving term $\zeta_{t}:=\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}+{\sum_{i=1}^{N}\delta_{i}}+\|\xi_{t}\|$ is uniformly bounded in expectation, i.e. there exists $C<\infty$ such that $\sup_{t}\mathbb{E}[\zeta_{t}]\leq C$, and that $\sup_{t}\mathbb{E}[\epsilon_{t}]\leq E<\infty$. Then the discounted expected return gap is bounded independently of the rollout horizon $H$:
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]=\mathcal{O}(1).$
Corollary 2 establishes that in the discounted setting with contractive dynamics ($L_{f}<1$), the expected return gap remains bounded by a constant that does not depend on the rollout horizon $H$. Intuitively, discounting attenuates the influence of long-term deviations, while contractive dynamics prevent the state deviation $d_{t}$ from growing unboundedly even when both observation and action perturbations are present. A key observation is that the nonstationarity term $\sum_{i=1}^{N}\delta_{i}$ enters the bound only through the uniform driving term $\zeta_{t}$; as long as these accumulated model updates remain bounded, the discounted error cannot compound over time. Consequently, the effect of both perturbations and update-induced drift is geometrically damped by the factor $\gamma L_{f}<1$, guaranteeing that the overall return difference remains $\mathcal{O}(1)$ even as $H\to\infty$. Practically, this indicates that stable long-horizon adaptation is achievable provided that policy updates are sufficiently smooth and that the system dynamics do not amplify perturbations.
@@ A.3 Proof of robustness theories
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\lambda\epsilon_{s}\right).$}
@@
From Lemma 1, the state deviation satisfies, for $d_{0}=0$,
$d_{t}\;\leq\;(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}})\,L_{f}\sum_{j=0}^{t-1}L_{f}^{\,j}=(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}})\,L_{f}\,\frac{L_{f}^{\,t}-1}{L_{f}-1}=\mathcal{O}(L_{f}^{\,t})\,(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}).$
At each step, the reward difference obeys the Lipschitz bound
$|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\;\leq\;L_{r}\big(d_{t}+\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}\big).$
Taking expectations on both sides and using the deterministic upper bound on $d_{t}$, we obtain
$\mathbb{E}\big[|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\big]\;\leq\;L_{r}\big(d_{t}+\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}\big).$
Summing over $t=1,\dots,H$ gives
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\;\leq\;\sum_{t=1}^{H}L_{r}\big(d_{t}+\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}\big).$
Finally substituting $d_{t}=\mathcal{O}(L_{f}^{\,t})(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}})$ and evaluating the geometric series yields
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\;\leq\;\mathcal{O}(HL_{r}L_{f}^{H})\,\big(\lambda\epsilon_{s}+\epsilon_{\mathrm{offline}}\big),$
which completes the proof. ∎
@@
Let $d_{t}=\|s_{t}-s_{t}^{*}\|$. Under $L_{f}$-Lipschitz dynamics and executed action $a_{t}=\pi_{t}(s_{t})+\xi_{t}$,
$d_{t+1}\leq L_{f}\big(d_{t}+\epsilon_{t}+\|\xi_{t}\|\big),$
with $\epsilon_{t}\leq\epsilon_{\mathrm{offline}}+{\sum_{i=1}^{N}\delta_{i}.}$
@@
Then the Lipschitz dynamics give:
$d_{t+1}=\|f(s_{t},a_{t})-f(s_{t}^{*},a_{t}^{*})\|\leq L_{f}(\|s_{t}-s_{t}^{*}\|+\|a_{t}-a_{t}^{*}\|)\leq L_{f}(d_{t}+\epsilon_{t}+\|\xi_{t}\|).$
@@
Assume actions are perturbed as $a_{t}=\pi_{t}(s_{t})+\xi_{t}$ with $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$, and VLA models satisfy $\|\pi_{i}-\pi_{i-1}\|_{\infty}\leq\delta_{i}$}. Then:
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\sigma\sqrt{d}\Big).$}
@@
Start from the per-step recurrence (Lemma 2):
$d_{t+1}\leq L_{f}\big(d_{t}+\epsilon_{t}+\|\xi_{t}\|\big),$
and recall the uniform bound on model mismatch $\epsilon_{t}\leq\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}$.
Unrolling the linear recurrence with $d_{0}=0$ gives
$d_{t}\leq\sum_{j=0}^{t-1}L_{f}^{\,t-j}(\epsilon_{j}+\|\xi_{j}\|)\leq\sum_{j=0}^{t-1}L_{f}^{\,t-j}\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\|\xi_{j}\|\Big).$
Take expectation on both sides and use linearity:
$\mathbb{E}[d_{t}]\leq\sum_{j=0}^{t-1}L_{f}^{\,t-j}\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\mathbb{E}\|\xi_{j}\|\Big).$
Now apply Lemma 3, which yields $\mathbb{E}\|\xi_{j}\|\leq\sigma\sqrt{d}$ for every $j$. Consequently,
$\mathbb{E}[d_{t}]\leq\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\sigma\sqrt{d}\Big)\sum_{j=0}^{t-1}L_{f}^{\,t-j}=\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\sigma\sqrt{d}\Big)\cdot\mathcal{O}(L_{f}^{t}),$
where the finite geometric sum produces a factor of order $\mathcal{O}(L_{f}^{t})$.
Because the reward is $L_{r}$-Lipschitz,
$|r(s_{t},a_{t})-r(s_{t}^{*},a_{t}^{*})|\leq L_{r}\big(d_{t}+\epsilon_{t}+\|\xi_{t}\|\big).$
Taking expectation and summing over $t=0,\dots,H-1$,
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]$ $\leq L_{r}\sum_{t=0}^{H-1}\mathbb{E}[d_{t}]+L_{r}\sum_{t=0}^{H-1}\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\mathbb{E}\|\xi_{t}\|\Big)$
$\leq L_{r}\sum_{t=0}^{H-1}\mathcal{O}(L_{f}^{t})\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\sigma\sqrt{d}\Big)+L_{r}H\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\sigma\sqrt{d}\Big).$
Collecting geometric-series terms gives the claimed form
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\Big(\epsilon_{\mathrm{offline}}+\sum_{i=1}^{N}\delta_{i}+\sigma\sqrt{d}\Big).$